OpenStudy (anonymous):

Substitution would be the simplest method of solution for which system? 5/3 x = 6y and 2x -3y = 8 y = 2x and x + y = 28 3x - 2y =18 and 4x + 5y =60

OpenStudy (anonymous):

2nd

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

2nd is the simplest one

OpenStudy (anonymous):

So whats the answer ? is it the 2nd one in order, orrr ??

OpenStudy (anonymous):

y=2x and x+y=28 x+(2x)=28 3x=28 x=28/3